# USD: Unsupervised Soft Contrastive Learning for Fault Detection in Multivariate Time Series

### v1

@@ IV-E Soft Contrastive Learning Loss with Data Augmentation
The conventional contrastive learning (CL) loss is structured around a single positive sample contrasted against multiple negatives. To refine this, we have restructured the CL loss into a more nuanced form that utilizes labels for positive and negative samples, denoted by $\mathcal{H}_{\text{c1,c2}}$. Detailed explanations of this transformation from Eq. (10) to Eq. (11) are available in Appendix A. Details of SCL loss.
$\mathcal{L}_{\text{CL}}(x^{\text{c1}},\{x^{\text{c2}}\}_{\text{c2}=1}^{N_{K}})$ (11)
$=$ $-\sum_{j=1}\{\mathcal{H}_{\text{c1,c2}}\log Q_{\text{c1,c2}}+(1-\mathcal{H}_{\text{c1,c2}})\log\dot{Q}_{\text{c1,c2}}\},$
where $\mathcal{H}_{\text{c1,c2}}$ indicates if samples $i$ and $j$ have been augmented from the same source. $\mathcal{H}_{\text{c1,c2}}=1$ signifies a positive pair $(x^{\text{c1}},x^{\text{c2}})$, and $\mathcal{H}_{\text{c1},\text{c2}}=0$ indicates a negative pair. The term $Q_{\text{c1},\text{c2}}=\exp(S(z^{\text{c1}},z^{\text{c2}}))$ represents the density ratio, as defined and computed by the backbone network.
To mitigate issues arising from view-noise, we introduce Soft Contrastive Learning (SCL). This approach moderates the impact of erroneous sample classifications by variably weighting different samples, calculated through similarity assessments by the network itself. The SCL loss function is defined as:
$\mathcal{L}_{\text{SCL}}(x^{\text{c1}},\{x^{\text{c2}}\}_{c2=1}^{N_{K}})$ (12)
$=-\sum_{j=1}\{P_{\text{c1,c2}}\log(Q_{\text{c1,c2}})+(1-P_{\text{c1,c2}})\log(1-Q_{\text{c1,c2}})\},$
where $P_{\text{c1,c2}}$ represents a softened version of $\mathcal{H}_{\text{c1,c2}}$, acting as weights for the samples, and $Q_{\text{c1,c2}}$ is the density ratio.
@@
$P_{\text{c1,c2}}=\left\{\begin{array}[]{lr}e^{\alpha}\kappa(y^{\text{c1}},y^{\text{c2}})&\text{if }\mathcal{H}_{\text{c1,c2}}=1\\ \kappa(y^{\text{c1}},y^{\text{c2}})&\text{otherwise}\end{array}\right.,Q_{\text{c1,c2}}=\kappa(z^{\text{c1}},z^{\text{c2}}).$ (13)
where, $\alpha$ is a hyper-parameter in the range [0,1], infusing prior knowledge of the data augmentation relationship $\mathcal{H}_{\text{c1,c2}}$ into model training. The kernel function $\kappa(\cdot)$ maps high-dimensional vectors to probability values. For this paper, we employ the t-distribution kernel function $\kappa^{\nu}(\cdot)$ due to its flexibility in adjusting the closeness of distribution in dimensionality reduction mappings.
$\kappa^{\nu}(z^{\text{c1}},z^{\text{c2}})=\frac{\Gamma\left(\frac{\nu+1}{2}\right)}{\sqrt{\nu\pi}\Gamma\left(\frac{\nu}{2}\right)}\left(1+\frac{\|z^{\text{c1}}-z^{\text{c2}}\|_{2}^{2}}{\nu}\right)^{-\frac{\nu+1}{2}},$ (14)
where $\Gamma(\cdot)$ is the Gamma function, and $\nu$ represents the degrees of freedom, which dictate the kernel’s shape. The SCL loss employs a softer optimization target compared to the rigid structure of traditional CL loss, enhancing resilience against noisy labels. A comprehensive discussion on the differences between these loss functions is provided in Appendix A.3 SCL is better than CL in view-noise, along with proof that SCL maintains a higher signal-to-noise ratio when confronted with view-noise.
IV-F Optimization Objectives
@@
The entire USD framework, encompassing the RNN, GNN, and FLOW models, is optimized jointly through Maximum Likelihood Estimation (MLE) to ensure effective anomaly detection. The joint optimization process is formulated as follows:
$L_{\text{USD}}=\mathcal{L}_{\text{MLM}}+\mathcal{L}_{\text{SCL}},$ (15)
@@
where $\mathcal{L}_{\text{MLM}}$ is the MLE loss, and $\mathcal{L}_{\text{SCL}}$ is the SCL loss. The MLE loss is designed to maximize the likelihood of observing the transformed data points $\hat{z}_{i}$ under the model, while the SCL loss is designed to enhance the model’s sensitivity to subtle differences between normal and abnormal patterns. By embedding these mathematical formulations into each step of the USD framework, we ensure a rigorous approach to modeling and detecting anomalies in multivariate time series data, setting a new standard for unsupervised fault detection in complex systems.
IV-G Pseudocode
Algorithm. 1 shows how to train our model and how to obtain the selected features.
Algorithm 1 The USD algorithm
Input: Train windows Data: $\mathcal{D_{\text{1}}}=\{\mathbf{x_{\text{i}}}^{c}\}_{i=1}^{M}$, Test windows Data: $\mathcal{D_{\text{2}}}=\{\mathbf{x_{\text{j}}}^{c}\}_{j=1}^{N}$, Learning rate: $\alpha$, Epochs: $E$, Batch size: $B$, Network: $R_{\theta},G_{\phi},H_{\omega},f_{\alpha}$, loss weights: $\lambda$,
Output: $\hat{z}^{c}$ , the prediction of USD
1:  Initialization;
2:  while $i=0$; $i<E$; $i$++ do
3:     while $b=0$; $b<[M/B]$; $b$++ do
4:        $x^{c}$leftarrow Sampling($\mathcal{D_{\text{1}}}$, b); # Sample a batch data
5:        $x_{\text{new}}^{c}$leftarrow Augment($x^{c}$ ) by Eq. (9); # Data augmentation
6:        $g^{c}=R_{\theta}(x^{c}+x_{\text{new}}^{c})$ by Eq. (3); # Select the features
7:        $y^{c}$leftarrow $G_{\phi}(g^{c})$ by Eq. (4);
8:        $z^{c}$leftarrow $H_{\omega}(y^{c})$ by Eq. (5);
9:        $\hat{z}^{c}$leftarrow $f_{\alpha}(z^{c})$ by Eq. (6); # Make density estimates
10:        $\mathcal{L}_{\text{MLP}}\!\leftarrow\!L_{\text{MLP}}(\hat{z}^{c})$ by Eq. (7); # Cal. Maximum Likelihood Estimation (MLE) Loss
11:        $\mathcal{L}_{\text{SCL}}\!\leftarrow\!L_{\text{SCL}}(z^{c},\hat{z}^{c})$ by Eq. (12); # Cal. Soft Contrastive Learning (SCL)
12:        $\mathcal{L}\!\leftarrow\!\mathcal{L}_{\text{MLP}}+\mathcal{L}_{\text{SCL}}$ by Eq. (15);
13:        $\theta,\!\leftarrow\!\theta-\alpha\frac{\partial\mathcal{L}}{\partial\theta}$, $\phi\!\leftarrow\!\phi-\alpha\frac{\partial\mathcal{L}}{\partial\phi}$, $\bm{w}\!\leftarrow\!\bm{w}-\alpha(\frac{\partial\mathcal{L}}{\partial\bm{w}})$ Update the parameters
14:     end while
15:     while $b=0$; $b<[N/B]$; $b$++ do
16:        $x^{c}$leftarrow Sampling($\mathcal{D_{\text{2}}}$, b); # Sample a batch data
17:        $g^{c}=R_{\theta}(x^{c})$ by Eq. (3); # Select the features
18:        $y^{c}$leftarrow $G_{\phi}(g^{c})$ by Eq. (4);
19:        $z^{c}$leftarrow $H_{\omega}(y^{c})$ by Eq. (5);
20:        $\hat{z}^{c}$leftarrow $f_{\alpha}(z^{c})$ by Eq. (6); # Make density estimates
21:     end while
22:  end while
Return: $\hat{z}^{c}$

### v2

@@ IV-D Multi-Manifold Loss with Data Augmentation
The conventional contrastive learning (CCL) loss is structured around a single positive sample contrasted against multiple negatives. To refine this, we have restructured the CCL loss into a more nuanced form that utilizes labels for positive and negative samples, denoted by $\mathcal{H}_{\text{c1,c2}}$. Detailed explanations of this transformation from Eq. (6) to Eq. (7) are available in supplementary material.A. The $\mathcal{L}_{\text{CCL}}(x^{\text{c1}},\{x^{\text{c2}}\}_{\text{c2}=1}^{N_{K}})=$
$-\sum_{j=1}\{\mathcal{H}_{\text{c1,c2}}\log Q_{\text{c1,c2}}+(1-\mathcal{H}_{\text{c1,c2}})\log\dot{Q}_{\text{c1,c2}}\},$ (7)
where $\mathcal{H}_{\text{c1,c2}}$ indicates if samples $c1$ and $c2$ have been augmented from the same source. $\mathcal{H}_{\text{c1,c2}}=1$ signifies a positive pair $(x^{\text{c1}},x^{\text{c2}})$, and $\mathcal{H}_{\text{c1},\text{c2}}=0$ indicates a negative pair. The term $Q_{\text{c1},\text{c2}}=\exp(S(z^{\text{c1}},z^{\text{c2}}))$ represents the density ratio, as defined and computed by the backbone network. To enhance robustness against view-level noise introduced by data augmentation, we propose the Multi-Manifold Loss (MML). Unlike conventional contrastive losses that treat pairwise labels as binary constants, MML introduces soft similarity-based weights, enabling the model to down-weight uncertain or noisy samples during training.
Given an anchor sample $x^{c_{1}}$ and a set of associated samples $\{x^{c_{2}}\}_{c_{2}=1}^{N_{K}}$, the MML loss is defined as, $\mathcal{L}_{\text{MML}}(x^{c_{1}},\{x^{c_{2}}\})=$
$-\sum_{c_{2}=1}^{N_{K}}\left[P_{c_{1},c_{2}}\log Q_{c_{1},c_{2}}+(1-P_{c_{1},c_{2}})\log(1-Q_{c_{1},c_{2}})\right],$ (8)
where $P_{c_{1},c_{2}}$ is the soft label (weight) reflecting the similarity in input space, and $Q_{c_{1},c_{2}}$ denotes the similarity in the contrastive latent space. They are defined as,
@@
$P_{c_{1},c_{2}}$ =$=\begin{cases}e^{\alpha}\cdot\kappa(y^{c_{1}},y^{c_{2}})&\text{if}\mathcal{H}_{c_{1},c_{2}}=1\\ \kappa(y^{c_{1}},y^{c_{2}})&\text{otherwise}\end{cases},\quad$ (9)
$Q_{c_{1},c_{2}}$ $=\kappa(z^{c_{1}},z^{c_{2}}),$
where $\mathcal{H}_{c_{1},c_{2}}\in\{0,1\}$ indicates whether $x^{c_{2}}$ is a positive pair of $x^{c_{1}}$, $\alpha\in[0,1]$ is a confidence prior that emphasizes positive pairs, and $\kappa(\cdot,\cdot)$ is a similarity kernel.
To improve robustness against augmentation noise and sample-level variability, we adopt a generalized Gaussian kernel $\kappa^{\beta}(\cdot,\cdot)$ as the similarity function in MML. Compared to the standard Gaussian kernel, the generalized Gaussian provides a tunable shape parameter that allows for heavier tails when needed, enabling the model to assign meaningful similarity scores even to moderately distant pairs. The kernel is defined as:
$\kappa^{\beta}(a,b)=\exp\left(-\left(\frac{\|a-b\|_{2}}{\sigma}\right)^{\beta}\right),$ (10)
where $\beta>0$ controls the shape of the decay (with $\beta=2$ reducing to a Gaussian kernel, and $\beta<2$ producing heavier tails), and $\sigma$ is a scale parameter. This formulation makes the loss function more tolerant to view-level noise and non-uniform sample structures, leading to smoother gradients and improved generalization under multi-manifold distributions. Unlike USCL [23], which relies on a fixed Gaussian assumption and sharp contrastive margins, our generalized kernel formulation offers adaptive flexibility across manifold structures and improves tolerance to cross-view variation. Detailed comparisons and analysis are provided in Supplementary Material A.
IV-E Optimization Objectives
@@
The entire DMTFD framework, encompassing the RNN, GNN, and FLOW models, is optimized jointly through Maximum Likelihood Estimation (MLE) to ensure effective anomaly detection. The joint optimization process is formulated as follows,
$L_{\text{DMTFD}}=\mathcal{L}_{\text{MLE}}+\mathcal{L}_{\text{MML}},$ (11)
@@
where $\mathcal{L}_{\text{MLE}}$ is the MLE loss, and $\mathcal{L}_{\text{MML}}$ is the MML loss. The MLE loss is designed to maximize the likelihood of observing the transformed data points $\hat{z}_{i}$ under the model, while the MML loss is designed to enhance the model’s sensitivity to subtle differences between normal and abnormal patterns. By embedding these mathematical formulations into each step of the DMTFD framework, we ensure a rigorous approach to modeling and detecting anomalies in multivariate time series data, setting a new standard for unsupervised fault detection in complex systems.
@@ V Experiments
Datasets information. We evaluate DMTFD on six widely used multivariate time series (MTS) fault detection datasets (Table VIII). SWaT [11] and WADI [2] are collected from water treatment and distribution testbeds, simulating industrial cyber-attacks with labeled anomalies. PSM [1] contains server metrics from eBay, used to detect system performance anomalies. MSL [16] provides telemetry from NASA’s Curiosity rover for space mission anomaly detection. SMD [27] is from a large-scale server farm and captures system behavior across multiple machines; we report average results and ensure test sets include anomalies via fixed random seeds. SMAP is a NASA dataset [16] that contains data from a satellite’s attitude control system, used for detecting anomalies in space missions. All datasets are standard in one-class classification (OCC) for time series anomaly detection.
TABLE I: Fault detection performance of Area Under the Receiver Operating Characteristic (AUROC) on five public datasets. We compare the performance of different methods using a consistent window size (DMTFD) with our optimal results obtained through varying window sizes (DMTFD*) to showcase the maximum potential. The best results is in bold.
SWaT WADI PSM MSL SMD SMAP
DROCC [12] ICML, 2020 72.6(±3.8) 75.6(±1.6) 74.3(±2.0) 53.4(±1.6) 76.7(±8.7) 70.1(±3.5) 70.4 9
DeepSAD [26] ICLR, 2020 75.4(±2.4) 85.4(±2.7) 73.2(±3.3) 61.6(±0.6) 85.9(±11.1) 72.8(±4.1) 77.4 8
USAD [4] KDD, 2020 78.8(±1.0) 86.1(±0.9) 78.0(±0.2) 57.0(±0.1) 86.9(±11.7) 74.5(±2.9) 76.9 7
GANF [9] ICLR, 2022 79.8(±0.7) 90.3(±1.0) 81.8(±1.5) 64.5(±1.9) 89.2(±7.8) 77.6(±3.2) 80.5 6
Autoformer [30] NeurIPS, 2022 81.2(±1.1) 89.5(±1.3) 83.7(±1.8) 66.0(±1.5) 89.8(±7.1) 78.3(±3.1) 81.4 5
MTGFlow [39] AAAI, 2023 84.8(±1.5) 91.9(±1.1) 85.7(±1.5) 67.2(±1.7) 91.3(±7.6) 80.2(±2.8) 83.5 3
MTGFlow_C [40] TKDE, 2024 83.1(±1.3) 91.8(±0.4) 87.1(±2.4) 68.2(±2.6) 91.6(±7.3) 81.1(±2.7) 84.8 2
USD [23] ICASSP, 2025 82.7(±1.0) 90.5(±1.2) 84.5(±1.3) 66.8(±2.0) 90.5(±6.8) 79.5(±3.2) 82.4 4
DMTFD OURS 90.5(±0.9) 94.3(±0.4) 89.2(±2.0) 75.0(±2.2) 95.0(±6.1) 85.7(±2.5) 88.3 1
DMTFD* OURS 92.5(±1.2) 95.2(±1.1) 89.2(±2.1) 76.5(±1.0) 95.0(±6.4) 87.4(±2.1) 89.3 1*
TABLE II: Anomaly detection performance of Precision-Recall(PR) on five public datasets. We compare the performance of different methods using a consistent window size (DMTFD) with our optimal results obtained through varying window sizes (DMTFD*) to showcase the maximum potential. The best results is in bold.
SWaT WADI PSM MSL SMD SMAP
DROCC [12] ICML, 2020 26.4 (±9.8) 16.7 (±7.1) 60.7 (±11.4) 13.2 (±0.9) 20.7 (±15.6) 31.2 (±5.3) 28.2 9
DeepSAD [26] ICLR, 2020 45.7 (±12.3) 23.1 (±6.3) 66.7 (±10.8) 26.3 (±1.7) 61.8 (±21.4) 45.3 (±8.4) 44.8 7
USAD [4] KDD, 2020 18.8 (±0.6) 19.8 (±0.5) 57.9 (±3.6) 31.3 (±0.0) 68.4 (±19.9) 42.5 (±9.5) 39.8 8
GANF [9] ICLR, 2022 21.6 (±1.8) 39.0 (±3.1) 73.8 (±4.7) 31.1 (±0.2) 64.4 (±21.4) 50.7 (±4.3) 46.8 6
Autoformer [30] NeurIPS, 2022 35.1 (±7.2) 39.3 (±3.9) 71.4 (±6.0) 29.5 (±2.5) 60.2 (±18.1) 48.2 (±8.6) 47.3 5
MTGFlow [39] AAAI, 2023 38.6 (±6.1) 42.2 (±4.9) 76.2 (±4.8) 31.1 (±2.6) 64.2 (±21.5) 52.6 (±6.3) 50.8 3
MTGFlow_C [40] TKDE, 2024 41.2 (±4.8) 42.4 (±5.2) 76.1 (±4.9) 31.2 (±2.8) 64.5 (±20.1) 53.1 (±8.9) 51.4 2
USCL [23] ICASSP, 2025 37.4 (±5.9) 40.5 (±4.1) 74.2 (±5.6) 30.7 (±2.4) 63.3 (±18.4) 51.8 (±6.1) 49.6 4
DMTFD OURS 53.6 (±10.1) 45.7 (±4.1) 83.2 (±4.9) 32.1 (±3.5) 73.8 (±16.7) 60.2 (±7.2) 58.1 1
DMTFD* OURS 59.3 (±6.5) 60.5 (±3.2) 83.2 (±3.9) 46.2 (±1.8) 73.8 (±13.7) 65.3 (±8.9) 64.7 1*
Implementation details. For all datasets, we set the window size to 60-80 and the stride size to 10-30. All experiments were run for 400-600 epochs and executed using PyTorch 2.2.1 on an NVIDIA RTX 3090 24GB GPU. Additional specific parameters can be found in Supplementary Material.
V-B Evaluation Metric and Baselines Methods.
Following prior work, DMTFD performs window-level anomaly detection, where a window is labeled anomalous if it contains any anomalous point. We evaluate performance using two metrics: AUROC, which measures overall discriminative ability across thresholds, and AUPRC, which is more informative under class imbalance, capturing the trade-off between precision and recall.
Baselines. We compare DMTFD with several state-of-the-art (SOTA) anomaly detection methods. DROCC [12] is a robust one-class classification method assuming locally linear manifolds to avoid representation collapse. DeepSAD [26] is a semi-supervised method that detects anomalies via entropy differences in latent distributions. USAD [4] employs adversarially trained autoencoders for unsupervised time-series anomaly detection. GANF [9] enhances normalizing flows using Bayesian networks to model inter-series dependencies. MTGFlow [39] and MTGFlow Cluster [40] utilize dynamic graph structure learning and entity-aware flows for fine-grained density estimation, with the latter introducing clustering for improved accuracy.
V-C Fault Detection Performance on Five Benchmark Dataset
First, we discuss the performance advantages of the DMTFD method by comparing results on five common benchmarks. We list the results for the AUROC and AUPRC metrics in Tables I and  II, where the numbers in parentheses represent the variance from five different seed experiments. In Table I and Table II, the results of DROCC, DeepSAD, USAD, DAGMM, GANF, and MTGFlow are from the paper [39] [^1] https://github.com/zqhang/MTGFLOW. The results of MTGFlow Cluster are from [40] [^2] https://github.com/zqhang/MTGFLOW_Cluster. We report on two variants of DMTFD: one (DMTFD) uses the same hyperparameters across all experiments, and the other (DMTFD*) involves hyperparameter tuning specific to each target dataset. The best results are highlighted in bold. To visualize the benefits of DMTFD performance, the ROC and PR curves on four dataset is showned in Fig. 3. In the ROC and PR curves, the DMTFD method consistently outperforms the baseline methods, demonstrating its robustness and effectiveness in anomaly detection tasks.
Performance improvement. In terms of performance improvements, as shown in Tables I and Tables II, the DMTFD method shows significant enhancements compared to state-of-the-art methods. We compare the performance of different methods using a consistent window size (DMTFD) with our optimal results obtained through varying window sizes (DMTFD*) to showcase the maximum potential. The best results is in bold. These improvements are evident across both PR and AUC curves, particularly on the SWAT dataset with a 5.4% increase in AUROC and a 7.7% increase in AUPRC, and on the MSL dataset with a 6% increase in AUROC and a 7.0% increase in AUPRC. On average, there is a 4.1% increase in AUROC performance and a 7.0% increase in AUPRC performance. The consistent improvements in both AUROC and AUPRC indicate that the DMTFD method has enhanced performance across multiple aspects.
Stability improvement. As shown in Tables I and  II and Fig. 3, the stability of the DMTFD method is also noteworthy, exhibiting lower variance across all test datasets compared to traditional methods. This indicates that the DMTFD method is more robust and less sensitive to random initialization, making it a reliable choice for anomaly detection tasks.
Performance potential. Furthermore, by comparing the results of DMTFD and DMTFD*, we find that conducting hyperparameter searches for each dataset can significantly boost model performance. This suggests that our method has good hyperparameter adaptability and considerable potential for further performance enhancements.
Figure 4: Violin plot of AUROC scores on different training splits. The x-axis represents the training split ratio, and the y-axis represents the AUROC scores. The DMTFD method consistently outperforms the baseline methods(MTGFLOW) across different training splits, demonstrating its robustness and effectiveness in anomaly detection tasks.
Additional results on difference train/validation splits. To further investigate the influence of anomaly contamination rates, we vary training splits to adjust anomalous contamination rates. For all the above-mentioned datasets, the training split increases from 60% to 80% with 5% stride. We present an average result over five runs in Fig. 4. Although the anomaly contamination ratio of training dataset rises, the anomaly detection performance of MTGFlow remains at a stable high level. This indicates that the proposed DMTFD method is robust to the anomaly contamination ratio of the training dataset.
V-D Visualization of multi-subclass data distributions
Figure 5: Visualization of multi-subclass data distributions. (a) The t-SNE visualization of embedding of DMTFD on SWAT datasets. The color shown the different states of the data. Multiple clusters can be seen in the representation, which is consistent with the multiple manifold assumption of this paper. The normal and abnormal states are not a single cluster, but can be divided into separate subclusters. (b) The heatmaps of individual data point, each row in the heatmap is the index number of the time and each column is the index number of the sensor. We selected four sub-clusters, and two samples from each cluster were randomly selected for presentation. The similarity of samples within sub-clusters and the difference of samples between sub-clusters are illustrated.
Figure 6: Point plot of the anomalies predict outputs of MTGFLOW and proposed DMTFD. The x-axis represents the anomaly index, and the y-axis represents the log-likelihood of the anomaly. Anomalous ground truths are marked by a red background. This indicates that DMTFD is able to detect anomalies more accurately and in a timely manner.
To further investigate the effectiveness of the DMTFD method, we visualize the multi-subclass data distributions on the SWAT dataset. As shown in Fig. 5, the t-SNE visualization of the embedding of DMTFD on SWAT datasets reveals multiple clusters, consistent with the multiple manifold assumption of this paper. The normal and abnormal states are not a single cluster but can be divided into separate subclusters. The heatmaps of individual data points further illustrate the similarity of samples within sub-clusters and the differences between samples in different sub-clusters. This visualization demonstrates the ability of the DMTFD method to capture the diverse patterns present in both normal and abnormal states, enhancing its anomaly detection capabilities.
In addition, we present the point plot of the anomalies predict outputs of MTGFLOW and the proposed DMTFD in Fig. 6. The x-axis represents the anomaly index, and the y-axis represents the log-likelihood of the anomaly. Anomalous ground truths are marked by a red background. This indicates that DMTFD is able to detect anomalies more accurately and in a timely manner, outperforming MTGFLOW in terms of anomaly detection performance.
Fig. 7 shows the bar plot of the frequency of normalized anomaly scores on baseline methods and DMTFD. The x-axis represents the anomaly scores, and the y-axis represents the frequency of each score. The normalized anomaly scores of DMTFD are significantly lower than those of GANF and MTGFlow, indicating that DMTFD is more effective at distinguishing between normal and abnormal states.
Figure 7: The bar plot of the frequency of normalized anomaly scores on baseline methods and DMTFD. The x-axis represents the anomaly scores, and the y-axis represents the frequency of each score. The normalized anomaly scores of DMTFD are significantly lower than those of GANF and MTGFlow, indicating that DMTFD is more effective at distinguishing between normal and abnormal states.
V-E Ablation Study: The Effect of SoftCL Loss.
TABLE III: Ablation study. AUROC comparison on four datasets (SWaT, WADI, PSM, and MSL). The best results are in bold. The results are averaged over 5 runs, and the standard deviation is shown in parentheses.
Method SWaT WADI PSM MSL
DMTFD 90.5(±0.9) 94.3(±0.4) 89.2(±2.4) 75.0(±2.0)
w. Gaussian 88.9(±1.1) 93.9(±1.7) 86.7(±2.5) 72.1(±0.7)
w/o. MML 84.3(±1.5) 91.7(±1.2) 85.5(±1.3) 70.7(±2.4)
w/o. CL 84.7(±1.3) 90.8(±1.4) 85.9(±1.6) 69.2(±2.1)
w/o. Aug
84.8(±1.5) 91.9(±1.1) 85.7(±1.5) 67.2(±1.7)
To assess the effectiveness of each component designed in our model, we conducted a series of ablation experiments. The results of these experiments are presented in Table III.
We performed controlled experiments to verify the necessity of the Soft Contrastive Learning (SoftCL) Loss in the UNDA settings across all four datasets. The performance of the proposed DMTFD method is denoted as DMTFD in the table. The variant w/o. SoftCL represents the model performance with the SoftCL loss component, $\mathcal{L}_{MML}(\mathbf{x}_{i}^{t},{y_{i}}^{t})$, removed from the overall loss function of DMTFD. The variant w. CL indicates the model performance when the SoftCL loss is replaced by a typical Contrastive Loss (CL), $\mathcal{L}_{\text{CCL}}$, as defined in Eq. (6). The results clearly demonstrate that the SoftCL Loss significantly outperforms the traditional CL loss. We attribute the inferior performance of $\mathcal{L}_{\text{CCL}}$ to its inability to adequately address the view-noise caused by domain bias.
V-F Hyperparameter Robustness
Hyperparameter Robustness: Window Size and Number of Blocks. The Table. XII presents an ablation study investigating the robustness of hyperparameters, focusing on the window size and number of blocks. Three distinct datasets, SWaT, WADI, and PSM, along with MSL, are evaluated using DMTFD. Our method showcases promising results, with consistently competitive AUROC scores across varying hyperparameter configurations. Notably, the standard deviations accompanying AUROC values indicate a high degree of stability, underscoring the reliability of our approach. Analysis of the table reveals that optimal configurations often coincide with larger window sizes and moderate block numbers, suggesting a preference for capturing broader temporal contexts while maintaining computational efficiency. Additionally, trends indicate that as the number of blocks increases, there’s a discernible improvement in performance, albeit with diminishing returns beyond a certain threshold. This observation highlights the importance of carefully balancing model complexity with computational resources. Overall, the DMTFD method demonstrates robustness and effectiveness in anomaly detection tasks.
Hyperparameter Robustness: Learning rate, $\nu$, and number of epoch. In order to further investigate the effectiveness of MTGFlow, we give a detailed analysis based on SWaT dataset. We conduct ablation studies on the learning rate, $\nu$, and the number of epochs. The results are shown in Table. IV, Table. VI, Table. V and Table. VII. The results show that the performance of MTGFlow is relatively stable across different learning rates, $\nu$, and the number of epochs. This indicates that MTGFlow is robust to hyperparameter changes and can achieve good performance with a wide range of hyperparameters.
TABLE IV: Parameters analysis: learning rates. Table of AUC scores (AUROC%) for different learning rates on the SWaT dataset, indicating DMTFD’s sensitivity to the learning rate.
lr 0.001 0.005 0.01 0.05 0.1 0.5
TABLE V: Parameters analysis: number of epochs. Table of AUC scores (AUROC%) for different epochs on the SWaT dataset, indicating DMTFD’s sensitivity to the number of epochs.
epoch 40 100 200 400 500 1000
TABLE VI: Parameters analysis: hyperparameter $\nu$. Table of (AUROC%) scores (AUROC%) for different $\nu$ on the SWaT dataset, indicating DMTFD’s sensitivity to the hyperparameter $\nu$.
$\nu$ 0.005 0.01 0.05 0.1
AUROC 88.4(±1.8) 90.2(±0.9) 90.5(±1.3) 90.6(±2.6)
TABLE VII: Parameters analysis: wight of negative (Ne) sample and positive (Po) sample in Eq.(LABEL:equ_Lscl). Table of AUC scores (AUROC%) for different topological loss ratios on the SwaT dataset.
ne po 0.5 1 2 3 4 5
VI Conclusion
In this work, we propose DMTFD, an unsupervised framework for fault detection in multivariate time series. To overcome the limitations of Gaussian assumptions, DMTFD models data with multi-Gaussian representations. A neighbor-based augmentation strategy is designed to generate positive pairs for learning, while a representation head enables optimization based on local streamform similarities. This enhances the separation between normal and anomalous states and broadens anomaly coverage. Experiments on standard benchmarks (e.g., SWaT, WADI) demonstrate that DMTFD achieves superior AUC and PR scores, with lower false alarm rates and improved detection accuracy. Visualizations also validate the effectiveness of the multi-Gaussian modeling. Although our current focus is unsupervised fault detection, DMTFD can be extended to Remaining Useful Life (RUL) prediction. Its pretrained embeddings provide strong priors for temporal degradation modeling, and with minimal finetuning, can support few-shot RUL forecasting. We leave multi-task extensions integrating fault detection and RUL estimation for future work.
Acknowledgments
This work was supported by ‘Pioneer’ and ‘Leading Goose’ R&D Program of Zhejiang (2024C01140), Key Research and Development Program of Hangzhou (2023SZD0073), Beijing Natural Science Foundation (L221013), and InnoHK program.
A. Details of Dataset